OpenStudy (anonymous):

What is the geometric mean between 2 and 50?

OpenStudy (saifoo.khan):

10

OpenStudy (anonymous):

there you go!

OpenStudy (saifoo.khan):

\[\sqrt{2 * 50}\]

OpenStudy (anonymous):

\[\sqrt{2\times 50}=\sqrt{100}=10\]